40m
QIL
Cryo_Lab
CTN
SUS_Lab
CAML
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m Log |
Not logged in |
 |
|
Sat Dec 8 04:12:38 2012, Den, Update, LSC, prcl
|
Sun Dec 9 01:21:51 2012, Den, Update, LSC, beam inside DRMI is clipping on PR3 Tip-Tilt 6x
|
Sun Dec 9 17:51:53 2012, rana, Update, LSC, beam inside DRMI is clipping on PR3 Tip-Tilt
|
Mon Dec 10 03:02:03 2012, Den, Update, LSC, beam inside DRMI is clipping on PR3 Tip-Tilt
|
Mon Dec 10 10:13:41 2012, Den, Update, LSC, beam inside DRMI is clipping on PR3 Tip-Tilt
|
Mon Dec 10 22:34:34 2012, Den, Update, LSC, beam inside DRMI is clipping on PR3 Tip-Tilt   
|
Thu Dec 13 03:20:48 2012, Den, Update, LSC, beam inside DRMI is clipping on PR3 Tip-Tilt  
|
|
Message ID: 7800
Entry time: Sat Dec 8 04:12:38 2012
Reply to this: 7801
|
Author: |
Den |
Type: |
Update |
Category: |
LSC |
Subject: |
prcl |
|
|
Today I wanted to check that AS and REFL beams are real and contain proper information about interferometer. For this I locked YARM using AS55_I and REFL11_I. Then I compared spectrum with POY11_I locking. Everything is the same. I've also adjusted phase rotations of AS55 (0.2 ->24) and REFL11 (-34.150 -> -43).
Then I've locked MICH and aligned EMTs such that ASDC was close to zero. Then I locked PRCL and aligned PRM. Power buildup was 50.

|